OpenStudy (anonymous):

solve indefinite integral (lnx)^-1

OpenStudy (anonymous):

\[\int\limits 1/lnx\]

OpenStudy (anonymous):

let u=ln(x)

OpenStudy (anonymous):

du/dx = 1/x dx= x du = e^u du

OpenStudy (anonymous):

so then you have integral e^u / u , which should be possible to do by parts, hopefully

OpenStudy (anonymous):

@elecengineer... its not possible to solve by parts .. the series will continue on and on and on ...

OpenStudy (anonymous):

\[\int\limits u^{-1} e^u du \] m = e^u , u^(-1) du = dn dm/du = e^u , n = -1/(u^2)

OpenStudy (anonymous):

i tried the same

OpenStudy (anonymous):

yeh I wouldnt worry bout it, I looked at it for a few minutes then went to wolframa

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=1%2Fln%28x%29+

OpenStudy (anonymous):

if wolframa cannot do it then theres a pretty good chance that it cant be done by hand

OpenStudy (anonymous):

wait isnt their a reduction formula for integrals of the form x^n e^x

OpenStudy (anonymous):

yeah, it cant be done with normal techniques: http://en.wikipedia.org/wiki/Logarithmic_integral

OpenStudy (anonymous):

or is that only for positive n

OpenStudy (anonymous):

i checked wolframa too it says Lix wat is Lix?

OpenStudy (anonymous):

what i linked to, the Logarithmic Integral